OpenStudy (anonymous):

If x2 + y2 = 100, and xy = 18, find the value of x - y. How do I solve this?

OpenStudy (lgbasallote):

@waterineyes

OpenStudy (anonymous):

\[(x-y)^2 = x^2 + y^2 - 2xy\]

OpenStudy (anonymous):

\[(x-y)^2 = 64 \implies (x-y) = \pm 8\]

OpenStudy (anonymous):